### v1

# Communication Characterization of AI Workloads for Large-scale Multi-chiplet Accelerators

Abstract
Next-generation artificial intelligence (AI) workloads are posing challenges of scalability and robustness in terms of execution time due to their intrinsic evolving data-intensive characteristics. In this paper, we aim to analyse the potential bottlenecks caused due to data movement characteristics of AI workloads on scale-out accelerator architectures composed of multiple chiplets. Our methodology captures the unicast and multicast communication traffic of a set of AI workloads and assesses aspects such as the time spent in such communications and the amount of multicast messages as a function of the number of employed chiplets. Our studies reveal that some AI workloads are potentially vulnerable to the dominant effects of communication, especially multicast traffic, which can become a performance bottleneck and limit their scalability. Workload profiling insights suggest to architect a flexible interconnect solution at chiplet level in order to improve the performance, efficiency and scalability of next-generation AI accelerators.
Index Terms:
AI Accelerators, Multi-Chiplet Accelerators, Communication, Network-on-Package
I Introduction
Artificial Intelligence (AI) applications have revolutionized multiple fields such as natural language processing, genomics, medical and health systems, graph analytics, and data analytics, and others [1] [2] [3]. However, the impressive feats that AI can achieve are often accompanied by very intense computational demands, which are saturating the boundaries of the traditional computing infrastructures [4]. footnotetext: Authors gratefully acknowledge funding from the European Commission through grant HORIZON-ERC-2021-101042080 (WINC).
To address this, specialised hardware (HW) accelerators are designed that aim at executing specific computational-intensive AI workloads in an efficient manner. GPUs can be considered as hardware accelerators for graphics, that have lately adapted to also better support AI workloads. However, the past decade has seen the emergence and confirmation of specific AI accelerators with workload-specific operations such as, accelerators with tensor or transformer cores as the cornerstone of faster and more efficient AI [5]. Generally speaking, the architecture of AI accelerators is comprised of an off-chip memory, a global shared memory, known as a global buffer, and an array of processing elements (PEs) connected via a Network-on-Chip (NoC) [6].
As modern AI models are evolving in size and diversity to accommodate multiple applications [7] or to implement input-dependent models such as Graph Neural Networks (GNNs) [8], versatile and high-performance accelerators are required to support their execution. This can be achieved with reconfigurable dataflows in FPGAs [9], flexible accelerator architectures [10, 11], or by resorting to less efficient but more general-purpose GPU and CPU architectures [12].
Another alternative compatible with the scaling and adaptation of AI accelerators is the use of chiplets. Indeed, chiplet technology is a promising enabler that provides a way to scale AI accelerators, by combining together multiple specialized (and potentially heterogeneous) AI accelerator chiplets in a single platform, as illustrated in Figure 1. These chiplets are interconnected among themselves and to memory via on-package links, typically through silicon interposers or organic substrates, in order to create a Network-on-Package (NoP) [13, 14, 15]. This has been proposed in multiple works, including SIMBA [16], WIENNA [17] and others.
Figure 1: An illustration of multi-chiplet architecture with 3x3 computing chiplets and 4 DRAM chiplets.
It is worth noting that AI accelerator chiplets can spend more than 90 percent of the total system energy on memory-bound tasks, to fetch data, as illustrated in [18]. This is due to not only the limited speed of memory modules, but more importantly, the relatively slow speed of chiplet-to-chiplet data transfers, which can dominate the computation energy due to having to traverse long interconnects. This issue is exacerbated by the need to use multicast communication in many of the dataflows that are used in AI accelerators [17, 19].
Despite the importance of chiplet-level communication in multi-chip AI accelerators, the communication traffic within such systems has not been explored in depth nor characterized. Traffic characterization and modeling has been well studied in CPU and GPU platforms for a variety of workloads [20, 21, 22], yet a similar analysis is missing in the proposed context.
The main contribution of this work is the profiling of workload-specific data movements across a set of popular AI workloads in multi-chiplet AI accelerators of increasing size. We augment Gemini [23] to register the communication packets and then parse the resulting traces to analyze their characteristics, e.g. number of messages, destinations per multicast, or number of NoP hops per message.
The rest of the paper is organised as follows. Section II presents the state of the art work in multi-chip AI accelerators and workload characterization. Our methodology is presented in Section III. Section IV presents the results obtained, followed by conclusion and future work presented in Section V.
II Related Works
II-A Scheduler/ Mapper for AI Accelerators
In the pioneering work of [16], authors developed and characterised a multi-chip AI inference accelerator with a scalable and hierarchical interconnect architecture on a package, called SIMBA. The objective was to enhance the energy efficiency and reduce accelerator’s inference latency by partitioning the non-uniform workload, considering communication-aware data placement, and implementing cross-layer pipelining. The mapper used in this work is Timeloop [24], and the cost of HW architecture is estimated with Accelergy [25].
The authors of [23] developed and characterised a multi-chip AI inference accelerator design framework, called GEMINI, that explores the design space to deliver architectures that, for a particular workload, minimise monetary cost and Enery-Delay Product (EDP). The mapper employed in this work is inter-layer pipelining using SET [26], and a customised cost model was leveraged to evaluate the architecture cost. These improvements significantly enhanced overall system performance compared to the baseline mapping variants.
In recent works, the authors of [7] developed a multi-chiplet-based multi-model AI inference accelerator that is scalable under heterogeneous traffic (i.e., data center and AR/VR) models with the objective to minimise EDP. The mapper used in this work is based on SET [26], and the hybrid cost model is customised using MAESTRO [27].
In all the cases shown above, communication is either coarsely classified depending on the chosen dataflow or indirectly used as a metric for optimization of mapping. However, none of the previous works have done an in-depth characterization of the communication workload across AI algorithms and architectures.
II-B Workload Parallelism
Work that considered to determine spatial and temporal tiling factors such as, PE dimensions, accumulator SRAM sizing, scratchpad SRAM sizing, and finds the best compute/buffer sizes using a mapping-first approach includes Timeloop [24], Sparseloop [28], COSA [29] and DOSA [30] to reduce memory access overheads and data replication. Works that exploit both intra- and layer loop ordering using inter-layer pipelining include TANGRAM [31], SET [26], GEMINI [23], and most recently SCAR [7] to improve data re-usability reduce pipelining delays. Again, in these cases, communication cost is used directly or indirectly as a metric for optimization of the mapping, but without providing a proper characterization of the different workloads.
II-C Communication Workload Analysis
In addition, there are a few existing works that characterise the communication in CPU workloads based on cycle-accurate simulation, obtaining spatiotemporal distributions or correlations between traffic flows, among others, for unicast and multicast communication [20, 21, 32]. Similar analyses have also been performed in GPU architectures [22]. Undoubtedly, all the above work outperformed the performance on workloads-chiplets mapping, however, there is still a gap left to fulfil characterising workloads from inter-chiplet multicast communication perspective. This characterisation has significant potential to assist mappers/schedulers to efficiently orchestrate and tailor data movement strategies according to the underlying communication-distance bottlenecks in executing the next-generation AI workloads. To the best of our knowledge, there is a critical need to characterise multicast data movement pattern in-order to obtain the key metrics in the realm of diverse AI workloads-chiplet mapping under scalability (i.e., increasing the size of chiplet array configurations).
III Characterisation Methodology
The main objective of this work is to perform a traffic analysis to obtain communication behaviours of various AI workloads, aiming to assess their scalability. To this end, we employ the methodology summarised in Figure 2, to capture the important communication-related characteristics (e.g., multicast data movement) of all benchmark AI workloads.
Figure 2: Methodology for the characterization of data movement of AI workloads in multi-chip architectures, based on the Gemini framework [23].
The communication workloads to be analysed depend on multiple factors such as the AI workload itself, the size of the architecture, or how the workload is mapped into the architecture. The mapping requires decisions on the loop execution order and also regarding dataflow dependencies. The former determines the order in which the layer is mapped and executed on architecture, while later determines the inter-dependency of data provisioning between layers (i.e., the movement of output activation and partial sums, between preceding and subsequent layers).
To abstract away from these decisions, here we use GEMINI [23] to determine the best mapping for a particular combination of AI workload and multi-chip accelerator configuration. We wrote C++ functional scripts to obtain a set of communication traces from Gemini, which are later parsed with the aim to extract profiles, which can further be processed and visualised using Python 3.10. We mainly focused to capture multicast communication metrics, due to the severity of their effects when scaled, across all workloads and configurations. In particular, the communication metrics extracted are:
  • Amount of time in clock cycles spent on performing the data movement by relevant compute/memory chiplets.
  • Number of unicast and multicast messages.
  • Number of chiplet-to-chiplet (NoP) hops per message.
  • Number of chiplets involved in a multicast message.
Table I presents the design space exploration (DSE) parameters to map 12 AI inference workloads on best HW architecture parameters obtained from GEMINI. We choose benchmark models that contain multi-branch classic residual (e.g. ResNet50, ResNet152, GoogleNet, Transformer (TF), TF Cell), and inception (e.g. iRES) structures with more intricate dependencies which are prevalent and widely used in various scenarios such as image classification and language processing. The bandwidths in the different communication levels, as well as the three multi-chiplet arrangements from 2 to 18 chiplets, are chosen based on GEMINI baselines. Regardless of the number of AI acceleration chiplets, the number of DRAM chiplets is set to four.
TABLE I: DSE Simulation Parameters
Number of chiplets (1x2), (3x3), (3x6)
NoP XY Mesh, 4 GB/s/side
NoC XY Mesh, 8 GB/s/port
DRAM 4 Chiplets, 16 GB/s/chiplet
Darknet19, DenseNet, GNMT, GoogleNet,
AI Workloads LSTM, ResNet101, ResNet101, ResNet152,
ResNet50, ResNext50, TF, TF Cell, iRES
IV Workload Characterisation Results
This section presents characterisation results and discusses the insights obtained by visualising the profiles of executing AI workloads on different architecture parameters of multi-chip accelerators.
IV-A Communication Source Analysis
Figure 3 reveals the sources of data movement by assessing the time spent by NoC, NoP, and DRAM across all workloads and averaged across system sizes. Notably, it can be observed that in most cases, the inter-chiplet data movement is the highest contribution among the three, especially for workloads such as iRES, LSTM, or TF.
Figure 3: Fraction of overall execution time (in clock cycles) spent by on-chip, chip-to-chip and chip-to-DRAM data movement related tasks for each AI workload across configurations.
This aspect is further analyzed in Figure 4, which shows the amount of time spent on data movement (i.e., communication time) contributed by each NoC, NoP and DRAM blocks for the different system sizes considered. In each workload execution the increase in chiplet configuration has a direct influence in increasing the NoP time and indirect affect on NoC time as chiplets are increased the computations cores are also increased handling computations more gracefully. However, it is evident that some AI workloads such as LSTM and GNMT are heavily bottlenecked by the inter-chiplet communication time. On the contrary, Transformer (TF) workload is more memory-bounded workload.
Figure 4: Amount of data movement time spent by NoP, NoC and off-chip DRAM across chiplet array configurations.
As a summary, Figure 5 shows. in the form of a box plot, the fraction of data movement time spent by NoP over the sum of NoC, NoP and DRAM time for all the workloads as a function of the number of chiplets in the system. It is evident from the plot that the amount of NoP time is directly proportional to the increase in the network size, reaching values well over 50% and even beyond 70%, which will eventually throttle the execution of the AI workload. Therefore, there is need to improve and design flexible data movement schemes in order to eliminate the communication bottleneck.
Figure 5: The fraction of time spent by NoP over the total communication time across all chiplet array configurations.
IV-B Multicast Traffic Analysis
Figure 6 shows the amount of multicast messages generated by each workload and the number of chiplets that are included in the destination set of each multicast message. First, it is evident form the plot that algorithms such as DenseNet or TF have a high number of multicast messages. Moreover, it is observed that the importance of muilticasts with two chiplet destinations is high. However, the amount of multicast messages with higher number of visited chiplets is very significant, especially in workloads such as LSTM or GNMT, and also considering the logarithmic nature of the Y axis scale.
Figure 6: Amount of multicast messages and size of their destinations sets (in number of chiplets) across all AI workloads and for three chiplet array configurations.
Finally, Figure 7 shows the distribution of unicast and multicast messages as a function of the number of NoP hops required in the 3x6 chiplet array. This figure offers multiple insights. On one hand, the notion of spatial locality is not common in multi-chip AI accelerators, as the increasing number of messages with high number of NoP hops suggest. On the other hand, even though the number of multicasts is significantly smaller than the number of unicasts, they are more prone to establishing long-range connections.
Figure 7: Number of unicast and multicast messages across all AI workloads for the 3x6 chiplet array configuration.
To summarise, workload-specific layers are consisted of data-dependent operators, with different compute and memory requirements, along with data movements requirements. As number of chiplets per array increases, the number of hops also increases in the network which increases the communication-specific latency in moving data from memory to compute units. Hence, the workload-specific multicast profiling has potential to identify the underlying communication-distance bottlenecks and alleviating them by improving the interconnect fabrics. Employing modular interconnect communication fabrics (i.e., using wireless [33] and/or optical-enabled interconnects [34]) either is one of the promising technology enablers that aim to tackle the communication overheads and computation-related power consumption challenges in AI accelerators besides expanding on-chip memory also known as processing in-memory and processing near-memory, and adopting vertical chip stacking technologies such as 3D stacked memory.
Hence, scaling-driven architecture, enabled by wireless communication or optical communication, has huge potential to address these challenging end-to-end latency constraints in over all system with reliability and stability. For instance, in wireless interconnect fabrics, latency due to communication overhead can be reduced by utilising the intrinsic broadcast nature of the wireless networks by switching to the wireless interface [33]. Moreover, priority-based packet transmission of certain critical messages over others can be exploited to alleviate specific communication bottlenecks that happen if the workload is unbalanced within the multi-chip accelerator architecture.
V Conclusion and Future Work
We presented a multicast characterisation tool to obtain and analyse the workload-specific multicast communication bottleneck of the scalable multi-chip AI accelerator design framework called Gemini, observing that communication latency of various AI workloads has potential to contribute to the overall execution time. One novel aspect of our analysis compared to the state of the art is that we are able to capture the AI workload-specific communication characteristics, which we show to scale proportionally to the number of chiplets in the system. We posit that this analysis and modeling approach is critical to take into account the relevant trade-offs in designing the architecture for large-scale AI accelerator systems. Specifically, addressing the potential communication-distance bottlenecks by supplementing in-package flexible interconnects using technology enablers such as wireless-empowered or optical-empowered interconnects in the large-scale multi-chip AI accelerators.